# Effective LoRA Adapter Routing using Task Representations

### v1

@@ 1 Introduction
Effectively using these adapter pools during inference introduces a critical bottleneck: adapter routing. Given a user query, how do we efficiently identify and select an appropriate subset of adapters? Existing approaches for LoRA adapter routing formulate this as a granular retrieval problem, attempting to map each individual query directly to a small number of adapters, typically between one and four. However, this strategy faces several limitations. First, existing methods either rely on a static pool of adapters (Huang et al., 2024; Wu et al., 2024) or require access to the original adapter training data to construct a searchable index (Zhao et al., 2024; Chronopoulou et al., 2023), conditions that rarely hold for dynamic, public adapter pools where training data is proprietary or withheld for privacy reasons. Second, methods that overcome this issue by directly routing based on adapter parameter values (Ostapenko et al., 2024) incur routing overheads that scale linearly with the number of adapters and layers, making them expensive for large libraries. Finally, formulating routing as a direct matching problem between queries and adapters fails to exploit the inherent structure of the problem: adapters are not granular experts for individual queries, but specialized modules designed to solve specific tasks.
@@
We implement LoRAuter and evaluate our framework on a mixed-task benchmark comprising diverse tasks and $48$ adapters (Zhao et al., 2024). Our results demonstrate that LoRAuter achieves 101.2101.2% of the performance of an Oracle task-aligned adapter in in-domain settings, effectively outperforming the upper bound of selecting the perfect adapter for a query. In out-of-domain scenarios, where the model must generalize to unseen tasks, LoRAuter outperforms the strongest baseline, LoraRetriever (Zhao et al., 2024), by $5.2$ percentage points. We then scale our system to a pool of 1500+ adapters fetched from the wild, and show that LoRAuter achieves competitive performance to the original $48$ adapters on the same mixed-task benchmark. Thus, LoRAuter can extract useful routing signals from large, noisy adapter collections, offering a scalable and robust solution for open-ended LoRA serving.
@@
  • Extensive evaluation: We conduct comprehensive evaluations across large adapter pools, multiple model sizes, and ablation settings, demonstrating the effectiveness of LoRAuter. We open source our code (see Appendix A).
@@ 2 Background and problem description
MoLE  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [ICLR24]}}$ (Wu et al., 2024) $\surd$ $\times$ $O(NL)$
LoraHub  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [COLM24]}}$ (Huang et al., 2024) $\surd$ $\times$ $O(1)$
AdapterSoup  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [FindingsEACL23]}}$ (Chronopoulou et al., 2023) $\times$ $\surd$ $O(1)$
LoraRetriever  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [FindingsACL24]}}$ (Zhao et al., 2024) $\times$ $\surd$ $O(N)$
Arrow $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{[ICML24]}}$ (Ostapenko et al., 2024) $\surd$ $\surd$ $O(NL)$
@@ 2.3 Problem formulation
We formalize the problem of LoRA adapter routing. We consider a pool of $N$ LoRA adapters, $\Phi=\{\phi_{1},\phi_{2},\ldots,\phi_{N}\}$, available for routing, where each adapter $\phi_{i}=(B_{i},A_{i})$. In practice, an adapter is not made of a single pair of matrices, but of a pair of matrices for every layer of the model that benefits from the adapter (Hu et al., 2022). To simplify notation, we consider these adapter layers to be implicit.
@@ 2.4 Shortcomings of existing approaches
Many prior methods violate one or more of these desiderata. Mixture-of-Experts (MoE)-style approaches such as MoLE (Wu et al., 2024) rely on learned gating networks trained with task-specific data, which must be retrained for each task and limits applicability in open-ended settings. Similarly, LoraHub (Huang et al., 2024) learns adapter composition weights using validation data, requiring task-specific supervision at deployment time.
Other approaches avoid training but still depend on adapter training data. AdapterSoup (Chronopoulou et al., 2023) and LoraRetriever (Zhao et al., 2024) embed adapters using training data and perform routing by matching queries to these representations. While effective, such methods are unsuitable for public or proprietary adapter pools where training data is unavailable. In contrast, spectral routing methods such as Arrow (Ostapenko et al., 2024) eliminate both training data and learned routing components by operating directly on adapter parameters, but incur higher adapter selection overheads due to layer-wise routing.
@@ 3.3 Query-based task retrieval (➏, ➐, ➑, ➒)
Similar to (Zhao et al., 2024), we set $I$ to the following instruction: “Represent the sentence for similar task retrieval”, allowing the embedding to capture task similarity.
@@ 3.4 Adapter composition ()
While prior methods (Zhao et al., 2024; Chronopoulou et al., 2023) treat all retrieved tasks as equally relevant by using uniform averaging for composition, we use an input-aware fusion mechanism that dynamically adjusts the contribution of each adapter based on the input-task similarity. Specifically, the fused adapter outputs $h^{\prime}$ are obtained through a weighted combination of the top-$K$ adapters in the output space:
@@
where $\sum_{i=1}^{K}w_{i}=1$ where the weights $w_{i}$ correspond to probabilities $p_{i}$. This formulation adaptively integrates information from task-relevant adapters, allowing the model to emphasize those most aligned with the input while minimizing interference from unrelated tasks. Furthermore, composition in the output space yields better performance than composition in the parameter space (Zhao et al., 2024), as LoRA adapters often exhibit limited alignment, which hinders effective parameter merging (Stoica et al., 2025).
@@ 4 Evaluation
We next evaluate the effectiveness of the LoRAuter framework using representative benchmarks and against four state-of-the-art routing baselines. Our evaluation proceeds as follows: we first report main performance results under non-OOD, OOD, and semi-OOD settings (see Figure 6, Appendix B), and study robustness when scaling to large adapter pools (§4.2); we then conduct two ablation studies to assess the contributions of retrieval and composition (§4.3); we examine performance when task labels are replaced with data-driven pseudo-tasks obtained via clustering (§4.4); and finally, we evaluate the efficiency of adapter selection using Successive Halving (§4.5).
@@ 4.2 Main results
OOD and non-OOD settings. Figure 2 presents the results for both the models (Llama2-7B & Llama2-13B) and settings. The corresponding detailed numerical results are included in Table 6 (Appendix C.3). Our proposed LoRAuter achieves the strongest overall performance across each configuration. Under non-OOD setting and Llama2-7B, it attains an average normalized score of 101.2%, slightly surpassing the oracle (100%). This shows that uncertainty-aware fusion of complementary adapters from related tasks can sometimes even outperform the adapter trained for a given task. Under the OOD setting and Llama2-7B, LoRAuter reaches 88.4%, outperforming LoraRetriever, the best baseline by +5.2 points. The same pattern holds for Llama2-13B. LoRAuter achieves 98.8% on non-OOD evaluation, again closely matching the oracle, and 86.8% in the OOD setting. Spectral routing methods (Arrow and SpectR) achieve relatively lower performance as they directly route based on parameter values, which may not carry sufficient routing signal. As with the 7B model, LoRAuter outperforms baselines in each setting, demonstrating that the advantages of uncertainty-aware composition persist even on larger models.
Semi-OOD setting. We introduce a semi-OOD setting that lies between the non-OOD and OOD regimes and reflects realistic deployment scenarios. In this setting, the adapter trained on the target task is unavailable, but a validation set from the target task is accessible for routing. Figure 3 presents the results, with detailed numerical values reported in Table 6 (Appendix C.3). With access to validation data, LoRAuter can more accurately identify relevant adapters using task representations. For LoraRetriever, we continue to report OOD performance, as it does not leverage validation data and relies solely on training data. We find that LoRAuter improves to 92.7% in the semi-OOD setting from 88.4% in the OOD setting on Llama2-7B. Similarly, it improves to 91.1% from 86.8% on Llama2-13B. On both models, this corresponds to gains exceeding 4 points and remains substantially higher than LoraRetriever.
@@
We study the robustness of LoRAuter under large and heterogeneous adapter pool fetched “from the wild”. We collect all adapters available for Llama2-7B with rank $\leq 64$ from HuggingFace, resulting in a pool of 1567 adapters. We run LoRAuter on this pool which now does not contain the 48 well curated adapters but only adapters fetched from the wild. In the OOD setting, LoRAuter impressively achieves an average normalized performance of $85.7\%$ compared to $88.4\%$ obtained using the 48 adapter pool. Similarly, in the semi-OOD setting, LoRAuter achieves $89.6\%$ compared to $92.7\%$ on the well curated pool. This result demonstrates the potential of LoRAuter to extract useful routing signals from large, noisy adapter collections with no information about their training data whatsoever. To the best of our knowledge, this is the first work to evaluate LoRA adapter routing using such large collections of publicly available adapters. In this setting, the baseline methods cannot be meaningfully evaluated, as they either require access to adapter training data or remain computationally expensive to scale to 1500+ adapters. The detailed numerical results are included in Table 7 (Appendix C.3); for completeness, we also report results for a combined adapter pool that augments the HuggingFace adapters with the original 48 curated adapters.
@@ 6 Conclusions
Appendix A Reproducibility statement
All datasets and LoRA adapters used in our experiments are publicly available. To ensure full reproducibility, we release our implementation together with all hyperparameters, evaluation scripts, and model outputs at https://anonymous.4open.science/r/LoRAuter-5B2C/README.md.
Appendix B Evaluation details
@@
B.1 Details of baseline methods
@@
(Huang et al., 2024). LoraHub is a task-specific adapter–merging method that learns task-specific fusion weights over a set of LoRA parameters via black-box optimization. The method outputs a single fused adapter per task and does not support input-adaptive routing. Following the recommendations in the original paper, we randomly sample 20 adapters from the full pool to constrain the optimization search space. For evaluation, each input is answered using the fused adapter optimized for its corresponding task (i.e., we select the task-specific fused adapter based on the ground-truth task label). LoRAHub was trained using the same validation set as in our method, and we fix the number of optimization iterations to 48, matching the total number of adapter evaluations performed by our approach.
@@
(Zhao et al., 2024). A retrieve-then-compose method that ranks LoRA adapters using similarity between input and adapter representations, then fuses the top-$k$ adapters uniformly. This method assumes the availability of supervised training data for each task and requires homogeneous adapter training conditions. Our method was evaluated using the exact same set of adapters and the same text encoder as employed in the original LoraRetriever paper.
@@
(Ostapenko et al., 2024). A spectral routing method that applies SVD to adapter representations and selects experts based on spectral affinity with the input. It is training-free but introduces additional per-token inference overhead proportional to the number of adapters. Following the recommendations of the original paper, we use a temperature of $1$ for converting spectral affinities into routing weights, and we retrieve $k=3$ adapters during routing for consistency with our other baselines.
@@
(Fleshman & Durme, 2025). A recently proposed spectral routing variant that aligns input and adapter representations through eigenspace projections. It avoids training but generally underperforms methods that incorporate task-level context. While the method supports softmax-based weighting of the selected adapters, in our evaluation we follow the configuration used in the original paper and apply uniform weights instead. Furthermore, we retrieve $k=4$ adapters during routing, as suggested in the original paper.
@@ B.2 Dataset details
Appendix C Additional results
@@
C.1 Selection-based methods
@@
C.2 LoraRetriever weighted fusion
@@
C.3 Main results per benchmark task
@@
C.4 Adapter performance on the validation set
@@ C.5 Successive halving
C.6 Choosing $K$ for Input-Aware Adapter Fusion
@@
Appendix D Full results

### v2

@@ 1 Introduction
Effectively using these adapter pools during inference introduces a critical bottleneck: adapter routing. Given a user query, how do we efficiently identify and select an appropriate subset of adapters? Existing approaches for LoRA adapter routing formulate this as a granular retrieval problem, attempting to map each individual query directly to a small number of adapters, typically between one and four. However, this strategy faces several limitations. First, existing methods either rely on a static pool of adapters Huang et al. (2024); Wu et al. (2024) or require access to the original adapter training data to construct a searchable index Zhao et al. (2024); Chronopoulou et al. (2023), conditions that rarely hold for dynamic, public adapter pools where training data is proprietary or withheld for privacy reasons. Second, methods that overcome this issue by directly routing based on adapter parameter values Ostapenko et al. (2024) incur routing overheads that scale linearly with the number of adapters and layers, making them expensive for large libraries. Finally, formulating routing as a direct matching problem between queries and adapters fails to exploit the inherent structure of the problem: adapters are not granular experts for individual queries, but specialized modules designed to solve specific tasks.
@@
We implement LoRAuter and evaluate our framework on a mixed-task benchmark comprising diverse tasks and $48$ adapters Zhao et al. (2024). Our results demonstrate that LoRAuter achieves 101.2101.2% of the performance of an Oracle task-aligned adapter in in-domain settings, effectively outperforming the upper bound of selecting the perfect adapter for a query. In out-of-domain scenarios, where the model must generalize to unseen tasks, LoRAuter outperforms the strongest baseline, LoraRetriever Zhao et al. (2024), by $5.2$ percentage points. We then scale our system to a pool of 1500+ adapters fetched from the wild, and show that LoRAuter achieves competitive performance to the original $48$ adapters on the same mixed-task benchmark. Thus, LoRAuter can extract useful routing signals from large, noisy adapter collections, offering a scalable and robust solution for open-ended LoRA serving.
@@
  • Extensive evaluation: We conduct comprehensive evaluations across large adapter pools, multiple model sizes, and ablation settings, demonstrating the effectiveness of LoRAuter.
@@ 2 Background and problem description
MoLE  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [ICLR24]}}$ Wu et al. (2024) $\surd$ $\times$ $O(NL)$
LoraHub  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [COLM24]}}$ Huang et al. (2024) $\surd$ $\times$ $O(1)$
AdapterSoup  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [FindingsEACL23]}}$ Chronopoulou et al. (2023) $\times$ $\surd$ $O(1)$
LoraRetriever  $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{ [FindingsACL24]}}$ Zhao et al. (2024) $\times$ $\surd$ $O(N)$
Arrow $_{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\text{[ICML24]}}$ Ostapenko et al. (2024) $\surd$ $\surd$ $O(NL)$
@@ 2.3 Problem formulation
We formalize the problem of LoRA adapter routing. We consider a pool of $N$ LoRA adapters, $\Phi=\{\phi_{1},\phi_{2},\ldots,\phi_{N}\}$, available for routing, where each adapter $\phi_{i}=(B_{i},A_{i})$. In practice, an adapter is not made of a single pair of matrices, but of a pair of matrices for every layer of the model that benefits from the adapter Hu et al. (2022). To simplify notation, we consider these adapter layers to be implicit.
@@ 2.4 Shortcomings of existing approaches
Many prior methods violate one or more of these desiderata. Mixture-of-Experts (MoE)-style approaches such as MoLE Wu et al. (2024) rely on learned gating networks trained with task-specific data, which must be retrained for each task and limits applicability in open-ended settings. Similarly, LoraHub Huang et al. (2024) learns adapter composition weights using validation data, requiring task-specific supervision at deployment time.
Other approaches avoid training but still depend on adapter training data. AdapterSoup Chronopoulou et al. (2023) and LoraRetriever Zhao et al. (2024) embed adapters using training data and perform routing by matching queries to these representations. While effective, such methods are unsuitable for public or proprietary adapter pools where training data is unavailable. In contrast, spectral routing methods such as Arrow Ostapenko et al. (2024) eliminate both training data and learned routing components by operating directly on adapter parameters, but incur higher adapter selection overheads due to layer-wise routing.
@@ 3.3 Query-based task retrieval (➏, ➐, ➑, ➒)
Similar to Zhao et al. (2024), we set $I$ to the following instruction: “Represent the sentence for similar task retrieval”, allowing the embedding to capture task similarity.
@@ 3.4 Adapter composition ()
While prior methods Zhao et al. (2024); Chronopoulou et al. (2023) treat all retrieved tasks as equally relevant by using uniform averaging for composition, we use an input-aware fusion mechanism that dynamically adjusts the contribution of each adapter based on the input-task similarity. Specifically, the fused adapter outputs $h^{\prime}$ are obtained through a weighted combination of the top-$K$ adapters in the output space:
@@
where $\sum_{i=1}^{K}w_{i}=1$ where the weights $w_{i}$ correspond to probabilities $p_{i}$. This formulation adaptively integrates information from task-relevant adapters, allowing the model to emphasize those most aligned with the input while minimizing interference from unrelated tasks. Furthermore, composition in the output space yields better performance than composition in the parameter space Zhao et al. (2024), as LoRA adapters often exhibit limited alignment, which hinders effective parameter merging Stoica et al. (2025).
@@ 4 Evaluation
We next evaluate the effectiveness of the LoRAuter framework using representative benchmarks and against four state-of-the-art routing baselines. Our evaluation proceeds as follows: we first report main performance results under non-OOD, OOD, and semi-OOD settings (see Figure 6, Appendix A), and study robustness when scaling to large adapter pools (§4.2); we then conduct two ablation studies to assess the contributions of retrieval and composition (§4.3); we examine performance when task labels are replaced with data-driven pseudo-tasks obtained via clustering (§4.4); and finally, we evaluate the efficiency of adapter selection using Successive Halving (§4.5).
@@ 4.2 Main results
OOD and non-OOD settings. Figure 2 presents the results for both the models (Llama2-7B & Llama2-13B) and settings. The corresponding detailed numerical results are included in Table 6 (Appendix B.3). Our proposed LoRAuter achieves the strongest overall performance across each configuration. Under non-OOD setting and Llama2-7B, it attains an average normalized score of 101.2%, slightly surpassing the oracle (100%). This shows that uncertainty-aware fusion of complementary adapters from related tasks can sometimes even outperform the adapter trained for a given task. Under the OOD setting and Llama2-7B, LoRAuter reaches 88.4%, outperforming LoraRetriever, the best baseline by +5.2 points. The same pattern holds for Llama2-13B. LoRAuter achieves 98.8% on non-OOD evaluation, again closely matching the oracle, and 86.8% in the OOD setting. Spectral routing methods (Arrow and SpectR) achieve relatively lower performance as they directly route based on parameter values, which may not carry sufficient routing signal. As with the 7B model, LoRAuter outperforms baselines in each setting, demonstrating that the advantages of uncertainty-aware composition persist even on larger models.
Semi-OOD setting. We introduce a semi-OOD setting that lies between the non-OOD and OOD regimes and reflects realistic deployment scenarios. In this setting, the adapter trained on the target task is unavailable, but a validation set from the target task is accessible for routing. Figure 3 presents the results, with detailed numerical values reported in Table 6 (Appendix B.3). With access to validation data, LoRAuter can more accurately identify relevant adapters using task representations. For LoraRetriever, we continue to report OOD performance, as it does not leverage validation data and relies solely on training data. We find that LoRAuter improves to 92.7% in the semi-OOD setting from 88.4% in the OOD setting on Llama2-7B. Similarly, it improves to 91.1% from 86.8% on Llama2-13B. On both models, this corresponds to gains exceeding 4 points and remains substantially higher than LoraRetriever.
@@
We study the robustness of LoRAuter under large and heterogeneous adapter pool fetched “from the wild”. We collect all adapters available for Llama2-7B with rank $\leq 64$ from HuggingFace, resulting in a pool of 1567 adapters. We run LoRAuter on this pool which now does not contain the 48 well curated adapters but only adapters fetched from the wild. In the OOD setting, LoRAuter impressively achieves an average normalized performance of $85.7\%$ compared to $88.4\%$ obtained using the 48 adapter pool. Similarly, in the semi-OOD setting, LoRAuter achieves $89.6\%$ compared to $92.7\%$ on the well curated pool. This result demonstrates the potential of LoRAuter to extract useful routing signals from large, noisy adapter collections with no information about their training data whatsoever. To the best of our knowledge, this is the first work to evaluate LoRA adapter routing using such large collections of publicly available adapters. In this setting, the baseline methods cannot be meaningfully evaluated, as they either require access to adapter training data or remain computationally expensive to scale to 1500+ adapters. The detailed numerical results are included in Table 7 (Appendix B.3); for completeness, we also report results for a combined adapter pool that augments the HuggingFace adapters with the original 48 curated adapters.
@@ 6 Conclusions
Appendix A Evaluation details
@@
A.1 Details of baseline methods
@@
Huang et al. (2024). LoraHub is a task-specific adapter–merging method that learns task-specific fusion weights over a set of LoRA parameters via black-box optimization. The method outputs a single fused adapter per task and does not support input-adaptive routing. Following the recommendations in the original paper, we randomly sample 20 adapters from the full pool to constrain the optimization search space. For evaluation, each input is answered using the fused adapter optimized for its corresponding task (i.e., we select the task-specific fused adapter based on the ground-truth task label). LoRAHub was trained using the same validation set as in our method, and we fix the number of optimization iterations to 48, matching the total number of adapter evaluations performed by our approach.
@@
Zhao et al. (2024). A retrieve-then-compose method that ranks LoRA adapters using similarity between input and adapter representations, then fuses the top-$k$ adapters uniformly. This method assumes the availability of supervised training data for each task and requires homogeneous adapter training conditions. Our method was evaluated using the exact same set of adapters and the same text encoder as employed in the original LoraRetriever paper.
@@
Ostapenko et al. (2024). A spectral routing method that applies SVD to adapter representations and selects experts based on spectral affinity with the input. It is training-free but introduces additional per-token inference overhead proportional to the number of adapters. Following the recommendations of the original paper, we use a temperature of $1$ for converting spectral affinities into routing weights, and we retrieve $k=3$ adapters during routing for consistency with our other baselines.
@@
Fleshman & Durme (2025). A recently proposed spectral routing variant that aligns input and adapter representations through eigenspace projections. It avoids training but generally underperforms methods that incorporate task-level context. While the method supports softmax-based weighting of the selected adapters, in our evaluation we follow the configuration used in the original paper and apply uniform weights instead. Furthermore, we retrieve $k=4$ adapters during routing, as suggested in the original paper.
@@ A.2 Dataset details
Appendix B Additional results
@@
B.1 Selection-based methods
@@
B.2 LoraRetriever weighted fusion
@@
B.3 Main results per benchmark task
@@
B.4 Adapter performance on the validation set
@@ B.5 Successive halving
B.6 Choosing $K$ for Input-Aware Adapter Fusion
@@
Appendix C Full results